the difference between two numbers is 16. Five times the smaller is the same as 8 less than twice the larger. Find the numbers
Answer by checkley77(12844)

X-Y=16 OR X=16+Y
5Y=2X-8
5Y=2(16+Y)-8
5Y=32+2Y-8
3Y=24
Y=24/3
Y=8 ANS. FOR THE SMALLER NUMBER.
X=16+8=24 ANS. FOR THE LARGER NUMBER.
PROOF:
5*8=2*24-8
40=48-8
40=40